ESPERANTOLOGIO/ESPERANTO STUDIES

EES

Redaktoro / Editor

Christer Kiselman, Uppsala University, Department of Mathematics, P. O. Box 480, SE-751 06 Uppsala, Sweden. ⟨kiselman@math.uu.se⟩

Redakta komitato / Editorial Board

Detlev Blanke, Otto-Nagel-Straße 110, DE-12683 Berlin, Germany.

   ⟨blanke.gil@berlin.snafu.de⟩

José Borges Neto, UFPR/Departamento de Lingüística, Rua Gen. Carneiro 460,

BR-80060-150 Curitiba, PR, Brasil. ⟨borges@coruja.humanas.ufpr.br⟩ Probal Dasgupta, University of Hyderabad, Centre for Applied Linguistics & Translation Studies,

P. O. Central University, IN-500046 Hyderabad, India. ⟨pdgalts@uohyd.ernet.in⟩ Michel Duc Goninaz, Val Saint-Georges, Chemin de l’Échelle, FR-13100 Aix-en-Provence, France Gotoo Hitosi, Tohoku University, Faculty of Arts and Letters, Department of Linguistics, Sendai,

   JP-981 Japan. ⟨gothit@sal.tohoku.ac.jp⟩

Edmund T. Grimley Evans, St. John’s College, Cambridge CB2 1TP, England.

   ⟨edmun.grimley-evans@cl.cam.ac.uk⟩

Koutny Ilona, Adam Mickiewicz Universitato, Lingvistika Instituto, Mie_{}dzychodzka 5,

PL-60-371 Poznań, Poland. ⟨ikoutny@amu.edu.pl⟩ Erich-Dieter Krause, Universität Leipzig, Ostasiatisches Institut, Schillerstraße 6, DE-04109 Leipzig,

Germany Sergej Kuznecov, Tverskaja 19-67, RU-103050 Moskva, Russian Federation.

   ⟨kuznecov@philol.msu.ru⟩

Jouko Lindstedt, Slavonic and Baltic Department, P. O. Box 4 (Vuorikatu 5 B), FI-00014 University

   of Helsinki, Finland. ⟨jouko.lindstedt@helsinki.fi⟩

Liu Haitao, Jisuanji Zhan, Qinghai Lu Chang, Xining, Qinghai, CN-810108, China.

   ⟨htliu@yeah.net⟩

Ingrid Maier, Uppsala University, Department of Slavic Languages, P. O. Box 527, SE-751 20 Uppsala,

   Sweden. ⟨ingrid.maier@slaviska.uu.se⟩

Geraldo Mattos, Rua João Motter, 98, BR-81210-260 Curitiba, PR, Brazil.

   ⟨mattos@super.com.br⟩

Marc van Oostendorp, Meertens Institut, Joan Muyskenweg 25, NL-1096 CJ Amsterdam, The

   Netherlands. ⟨marc@vanoostendorp.nl⟩

Sergio Pokrovskij, ul. Tereŝkova 28-32, RU-630090 Novosibirsk, Russian Federation.

   ⟨pok@iis.nsk.su⟩

Humphrey Tonkin, University of Hartford, Office of the President Emeritus, Mortensen Library,

West Hartford, CT 06117-1599, USA. ⟨tonkin@mail.hartford.edu⟩ John C. Wells, University College of London, Dept. of Phonetics and Linguistics, Gower Street,

London WC1E 6BT, England. ⟨wells@phonetics.ucl.ac.uk⟩ Yamasaki Seikô, JP-185-0014 Tôkyô, Kokubunzi, Higasi Koigakubo 3-18-12, Japan.

   ⟨syama@dokkyo.ac.jp⟩